### v1

@@ 4 Methods
For the long document summarization task, we consider Llama 3.1 (8B) [40], Mistral (7B) [41], Gemma1 (8B) [39], ChatGLM4 (9B) [38], ChatGLM_lora, Gemini 1.5 (Flash) [44], Claude 3.5 (Sonnet) [45], GPT-3.5 [42], and GPT-4 [43], as baselines. We do not require models’ ability of accepting images as inputs. The former five models are open-source LLMs, while the latter two models are closed-source FMs. Llama 3.1 (8B) is an open-source LLM proposed by Meta with diverse training datasets. Mistral (7B) is an open-source LLM proposed by Mistral. Gemma1 (8B) is an open-source LLM from Google. We use the instruct-tuning mode of these three models. ChatGLM4 (9B) is an open-source LLM proposed by ZhipuAI, and ChatGLM_lora is the fine-tuned version of this model using the training dataset from PubMed data. GPT-3.5, and GPT-4 are two closed-source models from OpenAI, and GPT-4 is more advanced than GPT-3.5.
For the scientific figure understanding task, we consider Qwen-VL-Chat [56], LLaVA-13B [57], Gemini 1.5 (Flash), GPT-4V, GPT-4o, o1, and Claude 3.5/3.7 (Sonnet) for evaluation. The former two methods are open-source, while the latter four models are closed-source. Qwen-VL-Chat is a large-scale LMM proposed by Alibaba Cloud. LLaVA-13B is an open-source LMM proposed by researchers from the University of Wisconsin–Madison and Microsoft Research. Gemini 1.5 (Flash) is a closed-source LMM proposed by Google. GPT-4V, GPT-4o, and o1 are three closed-source LMMs proposed by OpenAI. GPT-4o is a more advanced version than GPT-4V, while o1 provides stronger reasoning ability. We also consider the five major voting (o1 (5 mv)) baseline, which means we ask o1 five times and collect the outputs with the highest frequency as the final output; and best-of-5 oracle (o1 (bo5)), which means we determine if o1 can make correct answers among these five generated outputs. We also consider other reasoning models such as DeepSeek-R1 [71], but it does not support complicated figures as inputs. Claude is a closed-source LMM proposed by Anthropic.
@@
Codes Availability. Our codes can be found in: https://github.com/HelloWorldLTY/ARIEL. Our datasets can be found in https://huggingface.co/datasets/iLOVE2D/ARIEL_DATA and will be available after peer review. The license is an MIT license.
Device and Reproductivity. To reproduce our experiment results with open-source models, we utilize one NVIDIA H100 GPU for experiments. To reproduce our experiment results with closed-source models, we use official API and random state 2024 for querying, while other hyper parameters are kept as default settings. We have tested the randomness of generative models by including replicates from GPT-4 and o1, which is tolerant and will not affect our conclusion.
@@
This project is partly funded by OpenAI research program and Google Cloud research program.
@@ Appendix A Prompts
The COT prompt for text summarization is adding “Let’s think it step by step.” at the end of each input.
@@
3. Write a line of the form “The proposed solution is correct” or “The proposed solution is incorrect” at the end of your response based on your analysis.
@@ Appendix B Supplementary Figures
Extended Data Fig. 7: Boxplots with detailed metrics of comparing human assistants and LMM assistants for correcting answers.

### v2

@@ 4 Methods
For the long document summarization task, we consider Llama 3.1 (8B) (38), Mistral (7B) (39), Gemma1 (8B) (37), ChatGLM4 (9B) (36), ChatGLM_lora, Gemini-Flash-1.5 (45), Claude-Sonnet-3.5 (46), GPT-3.5 (42), GPT-4 (43), GPT-4o (44), GPT-OSS (20B) (41), GPT-OSS (120B) (41), and Deepseek-v3 (40) as baselines. We do not require models’ ability of accepting images as inputs. The former five models are open-source LLMs, while the latter two models are closed-source FMs. Llama 3.1 (8B) is an open-source LLM proposed by Meta with diverse training datasets. Mistral (7B) is an open-source LLM proposed by Mistral. Gemma1 (8B) is an open-source LLM from Google. We use the instruct-tuning mode of these three models. ChatGLM4 (9B) is an open-source LLM proposed by ZhipuAI, and ChatGLM_lora is the fine-tuned version of this model using the training dataset from PubMed data. GPT-3.5, GPT-4, and GPT-4o are three closed-source models from OpenAI, and GPT-4 (as well as GPT-4o) is more advanced than GPT-3.5. GPT-OSS series are open-source models from OpenAI. Deepseek-v3 is an advanced 670B large language model.
For the scientific figure understanding task, we consider Qwen-VL-Chat (57), LLaVA-13B (58), Gemini-Flash-1.5, GPT-4V, GPT-4o, o1, and Claude-Sonnet-3.5/3.7 for evaluation. The former two methods are open-source, while the latter four models are closed-source. Qwen-VL-Chat is a large-scale LMM proposed by Alibaba Cloud. LLaVA-13B is an open-source LMM proposed by researchers from the University of Wisconsin–Madison and Microsoft Research. Gemini-Flash-1.5 is a closed-source LMM proposed by Google. GPT-4V, GPT-4o, and o1 are three closed-source LMMs proposed by OpenAI. GPT-4o is a more advanced version than GPT-4V, while o1 provides stronger reasoning ability. We also consider the five major voting (o1 (5 mv)) baseline, which means we ask o1 five times and collect the outputs with the highest frequency as the final output; and best-of-5 oracle (o1 (bo5)), which means we determine if o1 can make correct answers among these five generated outputs. We also consider other reasoning models such as DeepSeek-R1 (74), but it does not support complicated figures as inputs. Claude is a closed-source LMM proposed by Anthropic.
@@
Codes Availability. Our codes can be found in: https://github.com/HelloWorldLTY/ARIEL, and the datasets are hosted on https://huggingface.co/datasets/iLOVE2D/ARIEL_DATA (to be released after peer review) under the MIT license. Experiments with open-source models were run on a single NVIDIA H100 GPU. To reproduce our experiment results with closed-source models, we use official API with random state 2024 and default parameters.
@@
This project is partly funded by OpenAI research program, Google Cloud research program, as well as NIH grants U01 HG013840, U24 HG012108, and P50 CA196530.
@@ Appendix A Prompts
The COT prompt for text summarization is adding “Let’s think it step by step." at the end of each input.
@@
3. Write a line of the form “The proposed solution is correct" or “The proposed solution is incorrect" at the end of your response based on your analysis.
@@ Appendix B Supplementary Figures
Extended Data Fig. 4: Case study based on the original abstract, human-generated abstract, and LLM-generated abstract.
@@
[…]